How would I solve tis equation by using the method of substitution.
-x+2y=2
x-2y=-2

Answer by ewatrrr(24785)

-x+2y=2 0r x =2y-2
x-2y=-2
(2y-2) - 2y = -2
-2 = -2 Infinite solutions. Same Line. 2y = x + 2